Sol Summary – December 7th

Crew 184 Daily Summary Report  06 Dec 2017

MDRS Daily Summary Report for Sol 5

Summary Title:  Back on schedule!

Mission Status:  The past two days have been hectic due to getting in to the swing of things on Mars and the numerous failures we have encountered.  As a consequence, our schedule fell by the wayside and we were largely winging it.  Yesterday we sat down, came to a stable configuration, discussed our lessons learned from our stay so far, and prepared to jump back onto our schedule.

We are still dealing with our power system failure, but today has been a day of focus where our team is efficiently moving from task to task and completing our assigned activities.

We’ve learned today that the low battery SOC we’ve been seeing since our mission started is likely due to a misconfigured battery panel.  Mission Control is seeking help to reprogram it and return our power system to nominal.  In the meantime we will continue to minimize power usage.  Due to SOC ~ 35% it is possible we lose power tonight, our team has prepared for this eventuality and everyone is prepared to handle the cold if necessary, and all team members have light sources on hand overnight.  We have also filled reserve containers with water due the expectation that if power is lost pipes will freeze overnight.

  1. Power system is still smart enough to accept power from generator once SOC starts dropping overnight.  System Nominal!
  2. Power system swaps to generator power.  Generator Failure overnight due to the cold.  Expected loss of power overnight
  3. Power System misconfiguration prevents accepting generator power.  Expected loss of power overnight.

In the meantime, activities!

This morning we have completed Anesthesia training for all crewmembers.  On a long duration mission it is critical crews know how to treat any injuries or other problems that arise, as part of this devising methods of training for crew on a long duration mission is critical.  Our Dr. has devised a mold that simulates arteries / nerves / and veins in a leg that allows crew to practice using ultrasound equipment to identify and isolate specific body parts.  With the training complete crew will be able to complete trial runs where we see how effective this is a training device and how crew performance increases over time.

We also completed our third Matryoshka EVA to SIte #3.  We continue to collect geological samples from our target sites, though this is the first day that our crew is starting to feel the fatigue from EVA, which is hard work!.  Today’s site was exceptionally rocky which gave a very ‘martian’ feel.

Our crew is looking forward to the weekend which we have schedule as ‘light duty’ days to allow ourselves to recuperate from a very busy week!

Sol Activity Summary:

1.      Matryoshka EVA Site #3

2.      Observatory Operation

3.      Anesthesia Training

4.      Reports and general HAB maintenance / checks

5.      Food inventory transfer to excel and submission to

6.      Exercise

7.  Brick Experiment

8.  Power System Monitoring

9.  Receipt and stowage of supply drop

10.  CPR techniques (ME-3)

11.  Media Visit Cancelled

Look Ahead Plan:

  1. Matryoshka EVA Site #4
  2. ANAM Winscat Baseline
  3. Musk Observation
  4. CPR Tecniques
  5. Brick Experiment
  6. Filming and Interviews
  7. Exercise / General HAB Maintenance

Anomalies in work:

1.      HAB Leak from under EVA door continuing (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.  Power System Misconfiguration

Weather: Still cold (below freezing).

Crew Physical Status: All crew in good physical health.

EVA: See above for EVA details

Reports to be filed:

Greenhab

Operations

EVA Request

Journalist

Astronomy

Sol Summary

Support Requested:

1.      Trash pickup from rear airlock (Non Burnable)

2.      Replacement ATV Fuel

Sol Summary – December 6th

Crew 184 Daily Summary Report  06 Dec 2017

MDRS Daily Summary Report for Sol 5

Summary Title:  Continuing failures, schedule replans, and a successful EVA!

Mission Status:  We are still waiting for a nominal morning!  To fix yesterday’s bandwidth problems we powered off all devices in the HAB, we then connected our ‘internet’ laptop to the modem and watched 100MB of bandwidth disappear in ~ 5 minutes.  We concluded there was a background process of some sort operating on that computer as no known programs were running.  We were able to fix this issue by swapping out this laptop and powering down the fault one.  Communication restored!

We exulted in this victory for approximately 10 minutes, then noticed that our generator had failed again and we were back to a powerdown configuration.  Our SOC was holding at 47%, but it was again clear that we were operating with a degraded power system.  Communication with Mission Control has tentatively identified the problem as due to a ‘summer fuel’ blend which has been noted before, which is separate from the oil leak we observed earlier.  The generator will likely continue to be faulty for the next few days until we can procure a ‘winter’ blend.  Winters on Mars can be tough!.  We are continuing to conserve power wherever possible and using the generator minimally to supplement the solar arrays when required.

We resumed nominal operations this afternoon with two primary activities

We decided to finish up the EVA that we left incomplete yesterday, going to Matryoshka Site #2 which is adjacent to the MDRS HAB.  Our EVA prep procedure continues to be challenging, but we are getting better.  We have allocated the best radios and headsets for EVA (recently delivered!) and placed them adjacent to the EVA prep room.  We have also pre-staged all EVA equipment (boots / gloves / first aid kid / emergency water supply / maps / etc . . . ) in the EVA room to minimize the chance that our prebreathe is interrupted by having to go out to the HAB to get supplies.  We have also started carrying backup radios with us since we have had to cut an EVA short due to radios running out of batteries.  The EVA training continues!

Once outside we successfully found samples from numerous geological strata in our target era and bagged them for later delivery and chemical analysis.  We had some excess time upon returning so we also took some of the PR images requested and will uplink them as part of the journalist report.

When the majority of the crew got back to the habitat we did an observatory training sessions where the crew astronomy gave a tutorial in observatory and telescope operation.  There’s something about being on Mars that gives observatory operations extra gravitas as you are already in the ‘space’ mindset.

Sol Activity Summary:

1.      Matryoshka EVA Site #2

2.      Observatory Training Session (no observing)

3.      HAB Cleaning

4.      Reports and general HAB maintenance / checks

5.      Food inventory transfer to excel

6.      Exercise

Look Ahead Plan:

  1. Matryoshka EVA Site #3
  2. Nerve Block Training
  3. CBS News Visit (unconfirmed – request confirmation on whether media crew is coming from Mission Support)
  4. Filming and Interviews
  5. Solar Observation
  6. Communication Experiment (Time Delay)
  7. CPR Technique Training
  8. Exercise / General HAB Maintenance

Anomalies in work:

1.      HAB Leak from under EVA door (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.      Generator failure

3.      Internet bandwidth problem was resolved today.

Weather: Still cold (below freezing), slightly windy

Crew Physical Status: All crew in good physical health.

EVA: See above for EVA details

Reports to be filed:

Greenhab

Operations

EVA Request

Journalist

Astronomy

Sol Summary

Support Requested:

1.      Trash pickup from rear airlock

2.      Replacement ATV Fuel

3.      Additional dish soap

4.  Confirmation of whether media visits will occur

Sol Summary – December 5th

Crew 184 Sol Summary Report  05 Dec 2017

MDRS Sol Summary Report for Sol 4

Summary Title:  Our first failure and major replan

Mission Status:  Today was a hectic day with three major failures our
team dealt with.
1.    Internet – Between 8:30am and 9:30am 800MB of bandwidth was
utilized. It is unknown what caused this but likely culprit was a
device being left on and doing a software update in the background.
Internet communication was lost at this time.  These reports are being
delivered by Shannon today.
2.    Water – During our morning fill of the loft tank the pump
was non-functional.  Possible causes were quickly narrowed down to
either a frozen pipe due to the below freezing temperatures or a
clogged filter. Our team immediately switched to using bottled water
which we keep as an emergency reserve while we troubleshot the issue.
At 1pm we called Mission Control for advice who recommended we connect
the pipe heating system and wait till midday, this resolved the issue
around 3pm and water was restored to the HAB
3.    Power – During our morning check of SOC it was noted that
the generator was not running, having failed sometime during the
night, and that SOC was at 3%.  We immediately instituted a shutdown
of MDRS powered equipment to preserve what little power we had and
investigated the generator.  It was noted that a significant oil leak
had developed from the oil cap, upon tightening the cap the leak
stopped and the generator was restarted.  When returning to the
science dome it was observed that SOC was hovering around 10% and did
not appear to be charging.  After consultation with Mission Control it
was discovered that the generator was in idle due to the power system
not requesting it to go to full delivered power given the fact that
the solar array system was charging the battiers (however slowly) .
In addition the SOC was likely incorrect due to it being a calculated
value and us confusing it due to going off scale.  We will assess SOC
tonight after a full day in the sun and decide forward plan.

Our morning was mostly consumed with maintenance and troubleshooting,
we quickly discovered that when receiving failures our level of
knowledge of the systems was exhausted in how to troubleshoot,
recommend future crew get access to and study manuals for MDRS
equipment prior to stay.

In the afternoon when we got our facilities in a stable configuratrion
four members of our team went on our first Matryoshka EVA. We were
able to go to the Kissing Camel ridge and collect two samples, however
we were unable to go to the other site in URC South due to delays in
getting out since we had to fix our non-nominal systems.

The two sample we collected were to highlight interesting
geomorphological features to be used in outreach activities back in
Oxford. The strike and dip of the ridge component the first sample was
extracted from was also taken. The intention of sample collection and
strike and dip measurement is to determine whether satellite and
outcrop scale imagery data is reliable at predicting these.

Sol Activity Summary:
1.    Troubleshooting and EVA as discussed above
2.    Exercise
3.    Assembling of exercise equipment using martian dirt
(medicine ball complete!)
4.    HAB Cleaning
5.    Solar observation

Look Ahead Plan:
With internet down we are unable to contact our scheduling team.  We
will receive the schedule in the morning when internet bandwidth
restarts and go over it as a team to plan our day.

Anomalies in work:
1.    Yamaha 250 ATV #3 wheel deemed non-recoverable.  We have
parked it off to the side for mission support.  Photos of damage
attached.

Weather: Very cold!  It was below freezing for most of the day.

Crew Physical Status: All in good health.

EVA: Delayed due to non-nominal systems, so we are requesting the same
EVA from Sol 3 but just for site A (mound on URC South) to finish off
the sample collection. The plan and crew utilised will be the exact
same as today.

Reports to be file:
Greenhab
Operations
Astronomy
EVA Request
Journalist
Sol Summary

Support Requested:
Trash placed in the Rear Airlock for pickup.  Burnable trash on the
side with the box.  Non-burnable trash on the other side.

Sol Summary – December 4th

MDRS Daily Summary Report for Sol 3

Summary Title: First full up day of activities!

Mission Status: Today really felt like our first day on mars. It was our first full activity day as well as our first day on EVA! Surprisingly it was consensus among our crew that our first full day was actually the easiest day we’ve had so far! I think the excitement of going outside and our team getting into the swing of things has really made us more efficient.

The primary activity for the day was an EVA on all crew to familiarize ourselves with EVA procedures, suit functionality, and operation of the rovers. Many lessons were learned, and it was good that we got the kinks out on a EVA solely suited to it. This will allow our scientific EVAs to be successful later on, as well as ensure that we can safely perform EVAs at a farther distance from the habitat.

Our team is following a specific protocol for depress / repress operations modeled after ISS EVAs. Our team is doing a 25 minute prebreathe in the EVA Prep room to prevent decompression sickness, once complete we are doing a 5 minute depressurization in the actual airlock. During repress we are allocating 5 minutes for repressurization options, though we have the capability to do an ’emergency repress’ if needed at the risk of rupturing crew eardrums.

Below are the several lessons learned we had during EVA.

  1. Peripheral vision is nearly non-existent when wearing the EVA helmets. This makes it difficult to travel together in an ATV convoy as it is impossible to turn around and check on the status of those behind you. We attempted to mitigate this by placing the radios in VOX and keeping in constant voice contact to maintain proper spacing, we also played ‘leapfrog’ with the ATV’s where the leader would stop after several minutes of driving to allow the two behind him to catch up and pass, ensuring no-one was left behind.
  2. We had several communication failures during EVA due to earpieces falling out of ears. We addressed this via using medical tape to secure the earpieces in place.
  3. We had to abort a prebreathe operation due to forgetting to bring supplies into the EVA Prep room. We instituted a supply check prior to closing the Prep Room door. If necessary to break EVA prebreathe protocol we do a 3 to 1 buy back, i.e. if you break prebreathe for 5 minutes, you add 15 minutes to prebreathe to compensate.
  4. Several of our crew members are trained scuba divers. We found that it was difficult to hear other EVA crew while driving the ATV due to engine and wind noise, we instituted a series of hand signals to allow non-verbal communication which quickly and concisely allowed EVA crew status to be conveyed.
  5. Do not wear high pony tails – recommend low braid. A high braid interferes with the way that your head can move within the EVA helmet and vision is confined to roughly a 45 degree angle down and lower.

During the high winds today we heard a bang as the outer airlock door was blown open. The inner airlock door held closed and thus we didn’t all die! We immediately closed the door to the EVA Prep room to form a second seal to the martian atmosphere. During tomorrows familiarization EVA we will enter the EVA Prep room, close the EVA Prep room door, and use that as a depressurization chamber. Immediately upon departure on tomorrows EVA we will repair the outer airlock door.

We also completed our first scientific objectives today, covered below.

Reduced Gravity CPR Experiment: When on Mars there is a possible complication when performing CPR. With your body weight reduced the typical CPR techniques are non functional due to reduced force on the recipient. Our team is practicing techniques under the supervision of our medical officer to allow force application in excess of body weight and measuring output using a springed scaled.

Communication Delay ‘Bricks’ Experiment: Our first two experiment runs of the brick experiment were performed. We started with a 0 time delay baseline, and increased to a 9 minute time delay with verbal information transfer (martian time delay). Our runs will continue to collect more data and determine the most efficient methods of information transfer.

We also completed some major maintenance tasks today

The Musk Observatory has been set up and put through it’s paces and is ready for operation.

Trailer Tank to Static Water Tank transfer operation – Crew completed the transfer this morning completely filling our static water tank. Heartfelt thanks to the Team supporting us for dropping the trailer tank and instructions on how to transfer the water. The MDRS Hab Manual was a roaring success in allowing us to complete this task!

Sol Activity Summary:

  1. Tasks mentioned above
  2. Exercise
  3. Dirt dehydration
  4. Science Dome organization and clean up
  5. Repair of Roof Hatch hole
  6. Our first shower! Heartily enjoying the new heater
  7. Completion of reports (soon)
  1. EVA Request
  2. Sol Summary
  3. Greenhab
  4. Journalist
  5. Operations
  6. Astronomer
  7. Inventory

WiFi range survey Radio check with all radios (all functional except for 1) Look Ahead Plan:

1. Matryoshka EVA #1

2. Communication ‘Lego Bricks’ experiment

3. Sleep Study / crew medical questionnaires

4. Musk Observatory Observation

5. Journalist interviews and filming

6. Medical status surveys with Crew Doctor throughout the day.

7. CPR Experiment

8. HAB Cleaning

9. Radio Range EVA – This activity is cancelled due to 1 EVA per day limit. It will be rescheduled for a future day. We will work with schedulers to ensure they understand this limit for the future.

Anomalies in work:

  1. ATV wobbly wheel
  2. Non Functional Power Strip (removed and replaced)
  3. 3 Non Functional Radios (does not impact operations)
  4. HAB paneling peeling off due to wind damag

Weather: Very mild weather. This was welcome after yesterday.

Crew Physical Status: All crew in good health.

EVA: 2 Familarization EVAs today

Support Requested:

  1. ATV Fuel
  2. Food Drop (not urgent)

.

Sol Summary – December 3rd

Crew 184 Daily Summary Report  03 Dec 2017

MDRS Daily Summary Report for sol 2

Summary Title:  Getting settled at MDRS!

Mission Status:  Today has been a busy day preparing both ourselves and the MDRS for the rest of our mission.  We spoiled ourselves with a hearty breakfast today (and to get rid of some of our fresh food!).

Today has been one of excitement preparing for our first EVA tomorrow.  We have encountered our first mission failure!

During the high winds today we heard a bang as the outer airlock door was blown open.  The inner airlock door held closed and thus we didn’t all die!  We immediately closed the door to the EVA Prep room to form a second seal to the martian atmosphere.  During tomorrows familiarization EVA we will enter the EVA Prep room, close the EVA Prep room door, and use that as a depressurization chamber.  Immediately upon departure on tomorrows EVA we will repair the outer airlock door.

We have discovered today that all the planning in the world still requires real-time flexibility as numerous activities took much longer than we expected.  For example the Exercise Bike took several hours and modifications to assemble and get working correctly, much more than we had allocated.  That being said, we have successfully completed all objectives and are ready to begin completing our science objectives tomorrow!.

Sol Activity Summary:

1.    Inventoried and organized the MDRS Hab and Greenhab

2.    Completed all reports

a.    Greenhab

b.    HSO

c.    Mission Plan

d.    Operations Report

e.    Sol Summary

f.     EVA Request

g.    Journalist

3.    Assembled exercise equipment (Stationary Bike / Weights)

4.    Got more detailed training from previous crewmembers

5.    Crew Exercise

6.    Internet troubleshooting (WiFi modem nonfunctional – connection working straight to modem)

7.    Cleaning and organizing MDRS

8.    Radio check with all radios (all functional except for 1)

9.    Communication check at all MDRS buildings with Habitat

10. Neuroblock Feasability Training

Look Ahead Plan:

1.    Familiarization EVA

2.    Communication ‘Lego Bricks’ experiment

3.    Sleep Study / crew medical questionnaires

4.    Musk Observatory Training and Observation

5.    Journalist interviews and filming

6.    Medical status surveys with Crew Doctor throughout the day.

7.    CPR Training

a.    Practice CPR techniques with crew medical doctor on CPR medical dummy

See schedule for tomorrow attached to this email.

Anomalies in work:

8.    Non Functional WiFi Modem

9.    Non Functional Power Strip (removed and replaced)

10. 1 Non Functional Radio

11. EVA Airlock Outer Hatch Failure (failed open) due to high winds.

12. Roof Hatch blown open continually (will close when winds die down)

Weather:  Very windy!  We have had two wind related MDRS failures.

Crew Physical Status:  All crew in good health.

EVA: No EVA today

Reports to be filed:

Greenhab

Operations

EVA Request

Journalist

Sol Summary

Mission Plan

Support Requested:

1.    We can’t find a template for Daily HSO Report.  Please provide what is desired in this report.

2.    Water Transfer from Delivery Tank to Static Tank

3.    More TB-Cide Quat Cleaner (we are out)

4.    Recommendation for wi-fi router failure.

a.    We have reset the router, checked our allocation, restarted laptop, and router is non-functional.  When hooking up laptop directly to modem internet works well.

Sol Summary – December 2nd

Sol 1

Summary: MDRS Training Day!

Author: Tom Horn

Mission Status:  Mission is starting great and we are enjoying our first day on Mars! Today was a busy day as we spent the day out of sim getting training from the previous crew and Shannon. We are excited for our first day ‘in sim’ tomorrow. We are taking a ‘light duty’ day tomorrow as we focus on finishing move in to the hab and getting acclimated. Due to this we are NOT doing an EVA tomorrow.

Notes for CAPCOM. Please communicate with our crew using this email address, mdrscrew184@gmail.com. This ensures our entire crew sees all communications. Please add this email address to the mission support distribution so in the future we can email these reports to you from that address.

One of the major research goals for our team is to follow a ‘Martian Day’ of 24:40 and study the effects on our team from following a longer day. Our team is going to follow the schedule below. We understand the the ‘CAPCOM Communication Window’ is between 7-9pm, we will always be awake during this time and will ensure to get reporting completed during this window, including any and all EVA requests. Note that we won’t be following a normal day so if a request is sent to our crew during our ‘sleep time’ our response will be delayed.

Further details will follow in the Mission Plan to be sent tomorrow.

Flight Day Calendar Day Wake Bed Twilight Start Sunrise Sunset Twilight End
Day 1 12/2 Not Scheduled 8:00 PM 6:53:57 AM 7:23:08 AM 5:02:24 PM 5:31:34 PM
Day 2 12/3 4:00 AM 8:40 PM 6:54:50 AM 7:24:03 AM 5:02:17 PM 5:31:30 PM
Day 3 12/4 4:40 AM 9:20 PM 6:55:41 AM 7:24:57 AM 5:02:13 PM 5:31:28 PM
Day 4 12/5 5:20 AM 10:00 PM 6:56:32 AM 7:25:50 AM 5:02:11 PM 5:31:28 PM
Day 5 12/6 6:00 AM 10:40 PM 6:57:21 AM 7:26:41 AM 5:02:10 PM 5:31:30 PM
Day 6 12/7 6:40 AM 11:20 PM 6:58:10 AM 7:27:32 AM 5:02:12 PM 5:31:34 PM
Day 7 12/8 7:20 AM 12:00 AM 6:58:57 AM 7:28:22 AM 5:02:16 PM 5:31:41 PM
Day 8 12/9 8:00 AM 12:40 AM 6:59:44 AM 7:29:10 AM 5:02:23 PM 5:31:49 PM
Day 9 12/10 8:40 AM 1:20 AM 7:00:29 AM 7:29:56 AM 5:02:31 PM 5:31:59 PM
Day 10 12/11 9:20 AM 2:00 AM 7:01:13 AM 7:30:42 AM 5:02:42 PM 5:32:11 PM
Day 11 12/12 10:00 AM 2:40 AM 7:01:55 AM 7:31:26 AM 5:02:54 PM 5:32:25 PM
Day 12 12/13 10:40 AM 3:20 AM 7:02:36 AM 7:32:08 AM 5:03:09 PM 5:32:41 PM
Day 13 12/14 11:20 AM 4:00 AM 7:03:16 AM 7:32:49 AM 5:03:26 PM 5:32:58 PM
Day 14 12/15 12:00 PM 4:40 AM 7:03:55 AM 7:33:29 AM 5:03:44 PM 5:33:18 PM
Day 15 12/16 12:40 PM Not Scheduled 7:04:32 AM 7:34:06 AM 5:04:05 PM 5:33:40 PM
Day 16 12/17 Not Scheduled – Handover Day 7:05:07 AM 7:34:42 AM 5:04:28 PM 5:34:03 PM

 

Sol Activity Summary: We arrived at 12pm today and spent the rest of the day learning about the MDRS with Shannon. We enjoyed our last ‘out of sim’ dinner at Stan’s burger shack with Shannon and the previous crew.

Look Ahead Plan:

Tomorrow our primary tasks are:

  1. Finishing move in to the HAB
  2. Completing an MDRS Inventory of supplies (food and other)
  3. Finish cleaning MDRS to start the mission.
  4. ScienceDome / Greenhab / Solar Observatory Detailed Inspection
  5. Completion of required reporting
    1. Mission Plan
    2. Health and Safety Officer Report
    3. If other ‘beginning of sim’ reports are requested, please let us know!

Anomalies in work:  None

Weather: Unseasonably warm!

Crew Physical Status: All crew in good physical health. Brandon Ferguson from the previous crew is staying with our crew until Monday night.

EVA: No EVA

Reports to be filed:

  1. Sol Summary (this is intended to include the commanders report as a narrative included in Mission Status. If this did not hit sufficient detail for this please let me know what details are desired and I will include them in future reports)
  2. Operational Report

Support Requested: None

Sol Summary – December 1st

Crew 183+ Sol Summary Report 1 December, 2017

Sol 4

Title: Curtis Arrives!

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

Today, we spent most of our time at the hab. At long last, Curtis arrives to fix the water heater in the hab, which has been out of commission for over a month now. He also insulated the water pipes and the outside tank. This morning, I was able to start my final batch of Sedimentation tests. All that’s left are my salinity & pH tests for all the soils.

Meanwhile, Chris offered to help put the rovers together. We now have three out of the four new rovers fully assembled.

At about 4:00 PM, I took a short EVA to Tank wash to resupply on a couple of samples that I needed.

Look Ahead Plan:

Tomorrow morning, Curtis will deliver and install Shannon’s water tank tomorrow, Shannon and I will be preparing the Lab and the Hab for Crew 184’s arrival

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy

EVA: Tank Wash area

Reports to be Filed:

-Operations Report

Support Requested: N/A

Sol Summary – November 30th

Crew 183+ Sol Summary Report 30November2017

Sol 3

Title: Further Exploration

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

Today, I spent from 10:00 AM to 2:00 PM, exploring the area south of Kissing Camel Range. I’d estimate that I covered anywhere from 6-8 km during that time frame. There were many interesting areas that I found, such as a couple of caves that a person can walk into. Although I didn’t find any dinosaur bones out there, I did find a lot of petrified wood. During this time, Shannon managed to complete the MDRS user manual that she’s been working on.

Between 3:00 PM & 5:00 PM, Shannon and I explored the last of the unchecked roads we still needed to check up on.

Look Ahead Plan:

Tomorrow, Curtis and his team will come and fix the water heater. Shannon and I will clean up the Lab for Crew 184.

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy

EVA: Region South of Kissing Camel Range & Road exploration

Reports to be Filed:

-Operations Report

Support Requested: N/A

Sol Summary – November 29th

Sol 2

Title: An Intriguing Find…

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

Today, I set off to explore the area to the South of Kissing Camel Range and West of Cow Dung Rd. First I investigated some of the washes crossing the road to see what I could find. In the first wash that I stopped in, I found pieces of petrified wood amongst other very interesting specimens. I’ll definitely have to explore this drainage basin later.

From there, I proceeded further south, until I reached a large wash near White Rock Canyon. I only walked about 2 minutes up, when I found a very large specimen that looked like mineralized bone marrow. Compared to the basalt ejecta littering the wash, this specimen looked completely different. The holes were all aligned in the same direction, and were about the same size. The basalt pieces had gas bubbles, but were of random size, with no obvious pattern. The size of the marrow was very large, with 3-4 mm spacing. If this was indeed a piece of a dinosaur bone, then that meant there is a very large dinosaur fossil somewhere in that drainage basin.

I left the specimen where I found it, planning to come back and visit later. After spending over 2 hours in this area, I didn’t find much more, though there is still quite a bit of ground to cover. I returned to the hab by around 2:00 PM, and informed Shannon on what I had found. On our way to town, we investigated the specimen that I had found in the wash. Shannon wasn’t sure what it was, but we confirmed that there weren’t any other rocks in that area that looked like it.  She figured that if it was indeed a remnant of a dinosaur bone, that it would have to come from a very large dinosaur, much larger than a Utahraptor or velociraptor. We plan on surveying this area further to see what we can find.

Soon after, we ran into town, found a socket wrench kit that would work with the bolts on the new rovers, and organized the next supply of food for Crew 184. We returned to hab after finishing our business in town.

Look Ahead Plan:

Tomorrow and the day after, I will be extensively searching the drainage basin where I found that specimen for large fossil exposures. Since Shannon told me that she hasn’t extensively explored this area, what I may find there could be very interesting.

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy

EVA:

Reports to be Filed:

-Operations Report

Support Requested: N/A

Sol Summary – November 28th

Crew 183+ Sol Summary Report 28November2017

Sol 1

Title: Departure of Crew 183

Written by: Brandon Ferguson

Mission Status: Good. Operations are Normal

Sol Activity Summary:

This was the last day for Crew 183. They packed the last of their belongings and headed off, back to Earth. After eating breakfast, Shannon and I started assembling the new rovers. It took over an hour to get just one close to being ready to go, as the rovers didn’t come with assembly instructions. Unfortunately, we don’t have a socket wrench that is compatible with the large bolts used with the rovers. Hopefully we’ll be able to find a socket large enough to work.

At around 2:30 PM, we headed off to Cactus Rd, to rediscover a road that was thought to be lost and forgotten to time. Our goal was to see if this area could be accessed and explored by future EVAs. It was obvious it hadn’t been used by ATVs or rovers in years, as the most prominent tracks were from cows. The cow tracks made the road extremely bumpy, slowing our progress significantly.

Once we reached the end, we found a wonderful place to access the Summerville Formation, an easy 100 meter walk. There were these red and yellow pieces of chert-like rocks scattered everywhere that could be petrified dinosaur poop.

Once we finished here, we made our way to the end of Yellow Zebra road and found an amazing place for future EVAs. This area could easily take several hours to explore adequately. As it got closer to dark, we made our way back to the Hab, and I resumed my soil analysis in the Lab.

Look Ahead Plan:

Tomorrow, we plan to gather food supplies for the next incoming crew. I also intend to start exploring the region south of Kissing Camel range, to see if I find anything interesting. There is also a road in the area we conducted our EVA today that we still need to explore, called Maxwell Montes Rd. We may either do this tomorrow or sometime before Saturday, when Crew 184 arrives.

Anomalies in work: No Problems

Weather: Clear, warm, calm winds

Crew Physical Status: OK, Healthy

EVA: End of Cactus Rd & Yellow Zebra Rd

Reports to be Filed:

-Operations Report

Support Requested: N/A